Reports 1-1 of 1 Clear search Modify search
Injection system (General activities)
menendez, chiummo, karathanasis - 16:30 Saturday 15 May 2021 (51803) Print this report
IMC baffle test

Today in collaboration with Antonino we studied the performance of the IMC baffle in different situations.

1º We switched on the baffle until the temperature of the ADCs reached a plateau. Once this point was reached the loops of the IMC were opened and corrections unplugged from the piscina. We let the IMC to cool down for ~25 min. This cooling down was also visible in the temperature sensors of the baffle (See Temperature_Holes.png from 11:00 to 11:30). Everything was restored back to the nominal condition at the end of this test.

2º We displaced the impact point of the beam in the mirror by modifying the offset of the quadrant photodiodes behind the IMC, this is also visible by looking at the spot on the mirror hitted by the beam using the infrared camera (e.g. Camera_103038_Diagonal.gif ). These movements had a clear effect on the sensors of the IMC as can be seen in figure Temperature_Holes.png or Sensors_2part.png in which we observe rises and falls in the number of counts corresponding to different impact positions. During one of the displacement high order modes appeared and remained during part of the test (See Camera_131859_Vertical_HM.gif). We restored everything back to the nominal condition once we finished our test (See 210515_instr_baffle_test_summary.png).

Analysis will follow.


Images attached to this report
Non-image files attached to this report
Comments to this report:
melo, chiummo - 14:12 Monday 17 May 2021 (51809) Print this report

Following the IMC baffle test, this morning we have analyzed the throughput losses of the cavity in two different moments: when the High Order Modes (HOM) have appeared and when the system was in normal conditions. 

On May 16 at 00:40h UTC ("normal" period), we had:

  •  mismatch = 7.4%
  • Output of IPC on EIB = 35.7 W (calculated with calibration of #50848)
  • IMC_TRA = 26.4 W;
  • "Other losses" = 0.5% (VIR-0225A-18)
  • In this conditions, we have 18.15% of throughput losses, corresponding to a round-trip loss of 605 ppm.

On May 15 at 13:18:49h UTC (HOM period), we had:

  • mismatch = 7.1%;
  • Output of IPC on EIB = 36 W;
  • IMC_TRA = 23 W;
  • "Other losses" = 0.5%;
  • Here, the calculated throughput loss is 28.5% (more than 10% compared to "normal" conditions), corresponding to a round-trip loss of 1048 ppm.
Search Help